• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 08:10:07 on October 9,2010

    Entry number 334028

    keyword=Owl shift summary

    We asked for restricted access at the beginning of the shift, as the target cooldown had finished. While lock-up was taking place, we took cosmics on the calorimeter to test new firmware.

    We got beam around 1:30, checked on Be0 and C-hole, with different raster sizes.

    We took a 1h-long run on the optics target at 10uA (when I tried to increase the current to a higher value, MCC realized there was an operational restriction to 1uA on the optics target).

    We moved to LH2 (LOOP 2, not Loop 1 as in the target GUI). Checked raster, and started taking data with and without sieve.

    Smooth data taking at the end of the shift.

    Runlist
    =========
    7064: Raster 2x2 on C-hole
    7066: Raster 5.5x4 on C-hole
    7067: Multifoil target
    7068: Raster 5.5x4 on LH2
    7069; Raster 7x7 on LH2
    7071: Sieve OUT, LH2, 10uA
    7072: Sieve IN, LH2, 10uA
    7073: Sieve IN, LH2, 10uA